1.) How is a common-size balance sheet created?

2.) Discuss how marketable securities are valued on the balance sheet.

3.) How can the allowance for doubtful accounts be used to assess earnings quality?

4.) Why is the valuation of inventories important in financial reporting?

5.) Why would a company switch to the LIFO method of inventory valuation in an inflationary period?

6.) Which inventory valuation method, FIFO or LIFO, will generally produce an ending inventory value on the balance sheet that is closest to current cost?

7.) Discuss the difference between the straight-line method of depreciation and the accelerated methods. Why do companies use different depreciation methods for tax reporting and financial reporting?

8.) What is the purpose of listing the account “Commitments and contingencies” on the balance sheet even through no dollar amounts appear?

9.) How is it possible for a company with positive retained earnings to be unable to pay a cash dividend?

10.) The King Corporation has total annual revenue of $800,000; expenses other than depreciation of $350,000; depreciation expense of $200,000 for tax purposes; and deprecation expense of $130,000 for reporting purposes. The tax rate is 34%. Calculate net income for reporting purposes and for tax purposes. What is the deferred tax liability?
Answer #1

Answer :

Question 1)

Common size Balance sheet :A common size balance sheet is a balance sheet that displays both the numeric value and relative percentage for total assets and total liabilities.In this balance sheet total asset is 100% and any asset line item is presented in relative percentage of total asset.

Example :

If total asset =$90

And cash =$27

Inventory =$9

Then in common size balance sheet

Cash =27/90=0.3=30%

Inventory =9/90=0.1=10%